LIFETIME ANALYSIS OF A DISSOCIATIVE STATE: OPTICAL-OPTICAL DOUBLE RESONANCE SPEC-TROSCOPY OF THE V=3 LEVEL OF THE A^{1}P STATE OF BH

\underline{A. GILKISON}, J. CLARK, C. R. VITERI and E. R. GRANT, Purdue University, Department of Chemistry, West Lafayette, IN USA..

The A ^{1}Π state of BH is characterized by its barrier to dissociation. We have observed new structure in the v=3, J=4-7 region near the top of the barrier by means of constant energy scans over intermediate states in double resonant transitions locked on single B^{1}Σ - X^{1}Σ transitions. Linewidths provide new information on the barrier height and tunneling lifetimes.